OpenStudy (anonymous):

i need to graph using the slope and y intercept with the following equation. I need points to graph x+5y=20

OpenStudy (anonymous):

re-write your eq as: y=-1/5 x+4 (slope intercept form)

OpenStudy (anonymous):

set x = 0, y = 4 (0,4) set y= 0, x = 20 (20,0) use those points

OpenStudy (anonymous):

crazy thing is my graph only goes to point 10

OpenStudy (anonymous):

In that case you can substitute any other x or y value to get a second point. i.e. set y = 5, x = -5 (-5,5) can be the second point